OpenStudy (aakashsudhakar):

I'm solving for a partial differential equation, and I'm getting caught up with the integration step. Can someone help? The integral is listed below.

OpenStudy (aakashsudhakar):

\[\int\limits_{0}^{\pi}(\sin x) ^2 (\sin nx) dx\]

OpenStudy (gorv):

\[\sin^{2}x = 1- \cos^{2}x\] \[\cos 2 x = 2*\cos^{2} x -1\]

OpenStudy (gorv):

\[\frac{ \cos2x +1 }{ 2 }= \cos^{2}x\]

OpenStudy (gorv):

\[\int\limits_{0}^{\pi}(1-\cos^2x)(\sin nx)*dx\]

OpenStudy (gorv):

\[\int\limits_{0}^{\pi}(1-\frac{ \cos2x+1 }{ 2 })*\sin n x * dx\]

OpenStudy (gorv):

\[\int\limits_{0}^{\pi}(1-\frac{ \cos2x }{ 2 }-\frac{ 1 }{ 2 }) *\sin nx * dx\]

OpenStudy (gorv):

solve it further and use sin C & D formula

OpenStudy (amistre64):

whats the partial diffy q it came from?

OpenStudy (irishboy123):

.

OpenStudy (irishboy123):

@AakashSudhakar this solves in another 3 or 4 lines: \[\int\limits_{0}^{π} (\frac{e^{ix} - e^{-ix}}{2i})^{2}(\frac{e^{inx} - e^{-inx}}{2i}) dx\]